OpenStudy (anonymous):

MEDALS FOR BEST ANSWERS. Solve for x: -3(x + 3) = -3(x + 1) - 5. Solve for x: 3x - 5 = 2x + 6. Solve for x: 3(x + 1) = -2(x - 1) - 4.

OpenStudy (anonymous):

-3(x + 3) = -3(x + 1) - 5 --- distribute through the parenthesis -3x - 9 = -3x - 3 - 5 --- combine like terms -3x - 9 = -3x - 8 -3x + 3x = - 8 + 9 0 = 1 (incorrect) There is no solution to this problem ================= 3x - 5 = 2x + 6 (subtract 2x from both sides) 3x - 2x - 5 = 6 (add 5 to both sides) 3x - 2x = 6 + 5 (combine like terms) x = 11 ================== 3(x + 1) = -2(x - 1) - 4 -- distribute through the parenthesis 3x + 3 = -2x + 2 - 4 -- combine like terms 3x + 3 = -2x + 2 -- add 2x to both sides 3x + 2x + 3 = 2 -- subtract 3 from both sides 3x + 2x = 2 - 3 -- combine like terms x = -1 any questions ?